The following sample observations were randomly selected:

X. 4. 5. 3. 6. 10

Y. 4. 6. 5. 7. 7

determine the correlation coefficient and the coefficient of determination



1
Expert's answer
2022-03-28T18:06:20-0400

x values

"\sum =28"

Mean 5.6

"\sum(x-\mu)^2=SS_z=29.2"

Y values

"\sum=29"

Mean =5.8

"\sum(y-\mu)^2=SS_z=6.8"

X and Y combined

N=5

"\sum (x-M_x)(y-M_y)=10.6"

r calculation:

"r=\sum (x-M_x)(y-M_y)/\sqrt{(SS_x)(SS_y)}=10.6/\sqrt{29.2)(6.8)}=0.7522"

Coefficient of determination

"r^2=0.7522^2=0.5658"